A local Computer City sells batteries ($2) and small boxes of pens ($6). In August, total sales were $972. Customers bought 6 times as many batteries as boxes of pens. How many of each did Computer City sell?
Answer by ikleyn(52794)

.
A local Computer City sells batteries ($2) and small boxes of pens ($6).
In August, total sales were $972. Customers bought 6 times as many batteries as boxes of pens.
How many of each did Computer City sell?
~~~~~~~~~~~~~~~

Let x = the number of boxes of pens;

then the number of batteries is 6x, according to the problem's description.


Next write the total money equation

    2*(6x) + 6x = 972  dollars.


Simplify and find x

    12x + 6x = 972

       18x   = 972

         x   = 972/18 = 54.


ANSWER.  54 boxes of penns and 6*54 = 324 batteries.

Solved.


Quite simple and straightforward; nothing complicated.